OpenStudy (anonymous):

HELP ME PLEASE! Show how to find the derivative of f(t)= t^2+1/t using the quotient rule. Then show how to find the derivative of the same function without using the quotient rule and first changing the form of the function. I get two different answers when I do this! HELP MEEEEEEEE

OpenStudy (anonymous):

the answer to the first part is \[2t + \frac{(t)(0) - 1(1)}{t^2}\] Simplify and you get \[2t - \frac{1}{t^2}\]

OpenStudy (anonymous):

its this all over t or just \[f(t)=t^2+\frac{1}{t}\]

OpenStudy (anonymous):

because if so you would not use the quotient rule the derivative of \[t^2\] is \[2t\] and the derivative of \[\frac{1}{t}\] is \[-\frac{1}{t^2}\]

OpenStudy (anonymous):

Now doing it without the quotient rule you first need to write 1/t in exponential form.. So the new function would look like this \[t^2 + t^{-1}\] Now use the power rule to find the derivative. \[2t + -1t^{-2}\] Simplify and you get \[2t - t^{-2}\] or \[2t - \frac{1}{t^2}\]. So you get the same answer.

OpenStudy (anonymous):

Thank you sooooooooooo much, it was on my take home test! :)